DEVELOPMENT OF A REDUCED-COST CHIRPED PULSE MICROWAVE SPECTROMETER

\underline{IAN A. FINNERAN}, DANIEL B. HOLLAND, P. BRANDON CARROLL, Department of Chemistry, California Institute of Technology, Pasadena, CA 91125; and GEOFFREY A. BLAKE, Divisions of Geological & Planetary Sciences and Chemistry & Chemical Engineering, California Institute of Technology, Pasadena, CA 91125.

Chirped pulse Fourier transform microwave (CP-FTMW) spectroscopy has become a ubiquitous technique in the highresolution molecular spectroscopy community. Unfortunately, many components of CP-FTMW spectrometers are extremely expensive. Here we report of the development of an inexpensive microwave circuit and we present spectra of tetrahydrofuran and methanol collected between 8-16 GHz. Possible applications in remote sensing will also be discussed.